# Flows of linear orders on sparse graphs

### v1

@@ 2.8. Reasonable expansions and the expansion property
Let $X(\mathcal{D})$ be the set consisting of the $L^{+}$-expansions $M^{+}$ of $M$ such that $M^{+}|_{A}\in\mathcal{D}$ for all $A\leq M$. (Here, $M^{+}|_{A}$ denotes the $L^{+}$-structure induced on the domain of $A$ by $M^{+}$.) We define a topology on $X(\mathcal{D})$: for $B\leq M$ with expansion $B^{+}\in\mathcal{D}$, we specify a basic open set $U(B^{+})=\{M^{+}\in X(\mathcal{D}):M^{+}|_{B}=B^{+}\}$. We will see below that $X(\mathcal{D})$ gives a $G$-flow.
Lemma 2.25.
Let $(\mathcal{K},\leq)$ be an amalgamation class with Fraïssé limit $M$. Let $\mathcal{D}$ be a reasonable $L^{+}$-expansion of $(\mathcal{K},\leq)$. Then for $A\leq M$ with expansion $A^{+}\in\mathcal{D}$, there is an expansion $M^{+}\in X(\mathcal{D})$ with $M^{+}|_{A}=A^{+}$.
Proof.
Write $M$ as the union of an increasing $\leq$-chain $A=A_{1}\leq A_{2}\leq\cdots$ starting at $A$. Let $A^{+}_{1}=A^{+}$. By part (4) of reasonableness, we may inductively expand $A_{i}$ to some $A^{+}_{i}$ ($i\geq 2$) such that $A^{+}_{i-1}\leq A^{+}_{i}\in\mathcal{D}$. We then take $M^{+}=\bigcup_{i\geq 1}A^{+}_{i}$. For $B\leq M$, we have that $B\leq A_{i}$ for some $i$, so by part (3) of reasonableness, $M^{+}|_{B}\in\mathcal{D}$. So $M^{+}\in X(\mathcal{D})$. ∎
Lemma 2.26 (after Prop. 5.3, [19]).
Let $\mathcal{D}$ be a reasonable expansion of the amalgamation class $(\mathcal{K},\leq)$. Let $M$ be the Fraïssé limit of $(\mathcal{K},\leq)$ and let $G=\operatorname{Aut}(M)$. Then $X(\mathcal{D})$ is a $G$-flow with the natural action.
Proof.
By Lemma 2.25, $X(\mathcal{D})$ is non-empty. It is straightforward to see that the action is continuous and $X(\mathcal{D})$ is Hausdorff. To see that $X(\mathcal{D})$ is compact, consider the topological space $Q=\prod_{A\leq M}\{A^{+}\mid A^{+}\text{ is an expansion of }A\,\wedge\,A^{+}\in\mathcal{D}\}$, where each set of the product has the discrete topology and $Q$ has the product topology. Using part (2) of reasonableness, we have that $Q$ is compact. We define a map $\gamma:X(\mathcal{D})\to Q$, $\gamma(M^{+})=(M^{+}|_{A})_{A\leq M}$, and it is straightforward to see that $\gamma$ is a homeomorphism onto its image. Thus it suffices to show that $\operatorname{Im}(\gamma)$ is a closed subspace of $Q$. We have that
$\operatorname{Im}(\gamma)=\{(p_{A})_{A\leq M}\in Q:\forall\,B\leq C\leq M,\,p_{C}|_{B}=p_{B}\},$
and so if an element of $Q$ lies in the complement of $\operatorname{Im}(\gamma)$, this is witnessed on a finite set. ∎
Definition 2.27.
@@
Let $(\mathcal{K},\leq)$ be an amalgamation class with Fraïssé limit $M$ and $G=\operatorname{Aut}(M)$. As the order expansion $\mathcal{K}^{\prec}$ is a reasonable expansion of $(\mathcal{K},\leq)$, we have that $X(\mathcal{K}^{\prec})$ is a $G$-flow: we denote this by $\mathcal{LO}(M)$, the flow of linear orders on $M$.
As $\mathcal{D}_{1}$ is a reasonable expansion of $(\mathcal{C}_{1},\leq_{1})$, we have that $X(\mathcal{D}_{1})$ is a $G_{1}$-flow, which we denote by $\operatorname{Or}(M_{1})$, the flow of orientations. (Recall that $M_{1}$ is the Fraïssé limit of $(\mathcal{C}_{1},\leq_{1})$ and $G_{1}=\operatorname{Aut}(M_{1})$.)
Lemma 2.28.
Let $\mathcal{D}$ be a reasonable expansion of the amalgamation class $(\mathcal{K},\leq)$. Let $M$ be the Fraïssé limit of $(\mathcal{K},\leq)$ and let $G=\operatorname{Aut}(M)$.
@@
Let $Y$ be a subflow of $X(\mathcal{D})$. Let $\mathcal{D}^{\prime}\subseteq\mathcal{D}$ be the class of finite $L^{+}$-structures which $\leq$-embed into some element of $Y$. Then
@@
(1): Parts (1), (2), (3) of reasonableness of $\mathcal{D}^{\prime}$ are straightforward, and (4) follows from the $\leq$-homogeneity of $M$ and the $G$-invariance of $Y$. Clearly $Y\subseteq X(\mathcal{D}^{\prime})$. To see that $X(\mathcal{D}^{\prime})\subseteq Y$, take $M^{+}\in X(\mathcal{D}^{\prime})$. It remains to show that $M^{+}\in\operatorname{cl}_{X(\mathcal{D}^{\prime})}(Y)=Y$: this follows from the $\leq$-homogeneity of $M$ and the $G$-invariance of $Y$.
(2): Let $A^{+}\in\mathcal{D}^{\prime}$. By the definition of $\mathcal{D}^{\prime}$ we may assume $A^{+}\leq M_{1}^{+}$ for some $M_{1}^{+}\in Y$. As $Y=\overline{G\cdot M_{0}^{+}}$, there is $g\in G$ such that $A^{+}=M_{1}^{+}|_{A}=(gM_{0}^{+})|_{A}$, so $A^{+}$leq-embeds into $M_{0}^{+}$ and thus $\mathcal{D}^{\prime}\subseteq\operatorname{Age}_{\leq}(M_{0}^{+})$. The reverse inclusion is immediate as $M_{0}^{+}\in Y$. ∎
Definition 2.29.
Let $\mathcal{D}$ be a reasonable expansion of the amalgamation class $(\mathcal{K},\leq)$. We say that $\mathcal{D}$ has the expansion property over $(\mathcal{K},\leq)$ if for $A\in\mathcal{K}$, there exists $B$ in $\mathcal{K}$ with $A\leq B$ such that for all expansions $A^{+},B^{+}$ of $A,B$ in $\mathcal{D}$, there exists a $\leq$-embedding $A^{+}\to B^{+}$.
It is easy to see that to prove the expansion property for $\mathcal{D}$, it suffices to show that for $A^{+}\in\mathcal{D}$, there exists $B\in\mathcal{K}$ such that for all expansions $B^{+}\in\mathcal{D}$ of $B$, there exists a $\leq$-embedding $A^{+}\to B^{+}$ (use weak coprecompactness of $\mathcal{D}$ and JEP for $\mathcal{K}$).
Theorem 2.30 ([17], Th. 4).
Let $\mathcal{D}$ be a reasonable expansion of the amalgamation class $(\mathcal{K},\leq)$ with Fraïssé limit $M$ and $G=\operatorname{Aut}(M)$.
Then the $G$-flow $X(\mathcal{D})$ is minimal iff $\mathcal{D}$ has the expansion property over $(\mathcal{K},\leq)$.
@@ 2.9. Meagre orbits
We now introduce the weak amalgamation property (WAP), which will be essential in the proof of Theorem 4.1. (WAP was first defined in [14]. A similar property (the almost amalgamation property) was previously defined in [12].)
Definition 2.31.
Let $\mathcal{D}$ be a reasonable class of $L^{+}$-expansions of an $L$-amalgamation class $(\mathcal{K},\leq)$. We say that $(\mathcal{D},\leq)$ has the weak amalgamation property (WAP) if:
for all $A\in\mathcal{D}$, there exists $B\in\mathcal{D}$ and a $\leq$-strong $L^{+}$-embedding $f:A\to B$ such that, for any $\leq$-strong $L^{+}$-embeddings $f_{i}:B\to C_{i}\in\mathcal{D}$ ($i=0,1$), there exists $D\in\mathcal{D}$ and $\leq$-strong $L^{+}$-embeddings $g_{i}:C_{i}\to D$ ($i=0,1$) with $g_{0}\circ f_{0}\circ f=g_{1}\circ f_{1}\circ f$. (Note that here we specify only that the diagram commutes for $A$.)
[…]
Lemma 2.32 ([4, Lemma 2.23] - adapting [14, Theorem 3.4]).
Let $\mathcal{D}$ be a reasonable class of $L^{+}$-expansions of an $L$-amalgamation class $(\mathcal{K},\leq)$. Let $M$ be the Fraïssé limit of $(\mathcal{K},\leq)$ and let $G=\operatorname{Aut}(M)$. Suppose that $X(\mathcal{D})$ is a minimal flow.
If $(\mathcal{D},\leq)$ does not have the weak amalgamation property, then all $G$-orbits on $X(\mathcal{D})$ are meagre.
For a proof, see [18, Lemma 1.77].
@@ 3. Admissible orders: a Ramsey expansion of $(\mathcal{D}_{1},\sqsubseteq_{s})$
We now provide an explicit description of a Ramsey expansion of $(\mathcal{D}_{1},\sqsubseteq_{s})$, given by the admissible orders on $(\mathcal{D}_{1},\sqsubseteq_{s})$, using Theorem 1.4 of [5]. This will be an essential tool in the proof of 4.1. (This Ramsey expansion will also have the expansion property over $(\mathcal{D}_{1},\sqsubseteq_{s})$, though we will not use this in the proof of Theorem 4.1.) In the below two definitions, we adapt definitions from [5, Section 1 and Section 3] to the specific case of $(\mathcal{D}_{1},\sqsubseteq_{s})$.
@@
(In the above, we adapt [5, Definition 3.5]. Several aspects of the general definition in [5] simplify in this case: (A4) can be omitted as closure components are single vertices, and (A6) follows from ()($\ast$)( ∗ ) and $(3)$.)
The below theorem is an immediate translation of [5, Theorem 1.4] to the context of this paper. We will explain how to adapt [5, Theorem 1.4] to our context at the end of this section.
@@
We now explain how to adapt [5, Theorem 1.4] and the definition of admissible orders found in [5] to give the definitions and theorem above. The paper [5] gives Ramsey expansions for classes of finite structures in languages that may include set-valued function symbols, which enables us to deal with the strong class $(\mathcal{D}_{1},\sqsubseteq_{s})$.
Definition 3.6 ([5, Section 1]).
@@
The above framework enables us to deal with $(\mathcal{D}_{1},\sqsubseteq_{s})$ as follows ([5, Section 5.1]). Let $\tilde{L}$ consist of the relational language $L_{\text{or}}$ of oriented graphs together with a unary set-valued function symbol $F$. Let $\tilde{\mathcal{D}}_{1}$ consist of the $\tilde{L}$ structures $\tilde{A}=(A,F_{A})$ where $A\in\mathcal{D}_{1}$ and $F_{A}:A\to\mathcal{P}(A)$ is a function sending each vertex of $A$ to its out-neighbourhood in $A$. Then there is a bijection $\mathcal{D}_{1}\to\tilde{\mathcal{D}}_{1}$ sending each $A\in\mathcal{D}_{1}$ to its unique $\tilde{L}$-expansion $\tilde{A}$ in $\tilde{\mathcal{D}}_{1}$, and for $A,B\in\mathcal{D}_{1}$, we have that $A\sqsubseteq_{s}B$ iff $\tilde{A}\subseteq\tilde{B}$. We then have that $\tilde{L}$-embeddings between elements of $\tilde{\mathcal{D}}_{1}$ are $\sqsubseteq_{s}$-embeddings when considered in the language $L_{\text{or}}$, and therefore $\tilde{\mathcal{D}}_{1}$ is a free amalgamation class.
We now recall [5, Theorem 1.4], which will give us an explicitly defined Ramsey expansion of $\tilde{\mathcal{D}}_{1}$ via admissibly ordered structures.
Theorem ([5, Theorem 1.4]).
@@
The above theorem, together with [5, Definition 3.5], which gives the explicit definition of admissible orders, gives a Ramsey expansion $\tilde{\mathcal{O}}_{1}$ of $\tilde{\mathcal{D}}_{1}$. Using the correspondence between $\tilde{\mathcal{D}}_{1}$ and $(\mathcal{D}_{1},\sqsubseteq_{s})$ detailed in the preceding paragraph, we thus obtain a Ramsey expansion $(\mathcal{O}_{1},\sqsubseteq_{s})$ of $(\mathcal{D}_{1},\sqsubseteq_{s})$ satisfying the conditions of 3.2 (this definition is just a direct adaptation of [5, Definition 3.5]).
@@ 4.1. Preparatory definitions and lemmas
Let $N_{1}=(M_{1},\rho)$ be the Fraïssé limit of $(\mathcal{D}_{1},\sqsubseteq_{s})$, and let $(N_{1},\prec_{\alpha})$ be the Fraïssé limit of $(\mathcal{O}_{1},\sqsubseteq_{s})$. (Here, we use 2.11: $(\mathcal{D}_{1},\sqsubseteq_{s})$ is a strong expansion of $(\mathcal{C}_{1},\leq_{1})$ and $(\mathcal{O}_{1},\sqsubseteq_{s})$ is a strong expansion of $(\mathcal{D}_{1},\sqsubseteq_{s})$.)
Recall that $G_{1}=\operatorname{Aut}(M_{1})$. Let $H_{1}=\operatorname{Aut}(N_{1},\alpha)$. By 3.3 and 2.20, we have that $H_{1}$ is extremely amenable.
@@ 4.2. Setup and proof notation
Let $Y\subseteq\mathcal{LO}(M_{1})$ be a minimal subflow of $G\curvearrowright\mathcal{LO}(M_{1})$. As $H$ is extremely amenable, the flow $H\curvearrowright Y$ has an $H$-fixed point $\prec_{\beta}$, and as $Y$ is a minimal $G$-flow, we have $Y=\overline{G\,\cdot\!\prec_{\beta}}$. Let $\mathcal{J}=\operatorname{Age}_{\leq_{1}}(M_{1},\prec_{\beta})$. By 2.28, we have $Y=X(\mathcal{J})$. We will show that $(\mathcal{J},\leq_{1})$ does not have the weak amalgamation property (WAP), which implies that all $G$-orbits on $Y$ are meagre by Lemma 2.32.

### v2

@@ 2.5. Flows from reasonable expansions
Let $L\subseteq L^{+}$ be relational languages. Let $(\mathcal{K},\leq)$ be an amalgamation class of $L$-structures with Fraïssé limit $M$, and let $\mathcal{D}$ be a reasonable $L^{+}$-expansion of $(\mathcal{K},\leq)$. Recall ([5, Section 2.3, Theorem 2.15]) that we obtain an $\operatorname{Aut}(M)$-flow $X(\mathcal{D})$ from $\mathcal{D}$ by taking $X(\mathcal{D})$ to be the set consisting of the $L^{+}$-expansions $M^{+}$ of $M$ such that $M^{+}|_{A}\in\mathcal{D}$ for all $A\leq M$. (here, $M^{+}|_{A}$ denotes the $L^{+}$-structure induced on the domain of $A$ by $M^{+}$), where the topology on $X(\mathcal{D})$ is given by: for $B\leq M$ with expansion $B^{+}\in\mathcal{D}$, we specify a basic open set $U(B^{+})=\{M^{+}\in X(\mathcal{D}):M^{+}|_{B}=B^{+}\}$.
@@
As the order expansion $\mathcal{K}^{\prec}$ of an amalgamation class $(\mathcal{K},\leq)$ (with Fraïssé limit $M$) is reasonable, we have that $X(\mathcal{K}^{\prec})$ is an $\operatorname{Aut}(M)$-flow: we denote this by $\mathcal{LO}(M)$, the flow of linear orders on $M$. As $\mathcal{D}_{1}$ is a reasonable expansion of $(\mathcal{C}_{1},\leq_{1})$, we have that $X(\mathcal{D}_{1})$ is a $G_{1}$-flow, which we denote by $\operatorname{Or}(M_{1})$, the flow of orientations.
We will need a very mild reformulation of [5, Lemma 2.16], a technical result regarding subflows of $X(\mathcal{D})$:
Fact 2.1.
@@
Let $\mathcal{D}$ be a reasonable expansion of an amalgamation class $(\mathcal{K},\leq)$ with Fraïssé limit $M$. Let $Y$ be a subflow of $X(\mathcal{D})$. Let $\mathcal{D}^{\prime}\subseteq\mathcal{D}$ be the class of finite $L^{+}$-structures which $\leq$-embed into some element of $Y$. Then:
@@
(1): [5, Lemma 2.16]. (2): Let $A^{+}\in\mathcal{D}^{\prime}$. We may assume $A^{+}\leq M_{1}^{+}$ for some $M_{1}^{+}\in Y$. As $Y=\overline{G\cdot M_{0}^{+}}$, there is $g\in\operatorname{Aut}(M)$ such that $A^{+}=M_{1}^{+}|_{A}=(gM_{0}^{+})|_{A}$, so $A^{+}$leq-embeds into $M_{0}^{+}$ and thus $\mathcal{D}^{\prime}\subseteq\operatorname{Age}_{\leq}(M_{0}^{+})$. The reverse inclusion is immediate as $M_{0}^{+}\in Y$. ∎
2.6. The expansion property
Let $(\mathcal{K},\leq)$ be an amalgamation class with Fraïssé limit $M$ and let $\mathcal{D}$ be a reasonable expansion of $(\mathcal{K},\leq)$. Recall that $\mathcal{D}$ has the expansion property over $(\mathcal{K},\leq)$ if for $A\in\mathcal{K}$, there exists $B$ in $\mathcal{K}$ with $A\leq B$ such that for all expansions $A^{+},B^{+}$ of $A,B$ in $\mathcal{D}$, there exists a $\leq$-embedding $A^{+}\to B^{+}$. Also recall the following fact:
Fact 2.2 ([5, Theorem 2.18]).
The $G$-flow $X(\mathcal{D})$ is minimal iff $\mathcal{D}$ has the expansion property over $(\mathcal{K},\leq)$.
@@ 2.7. Meagre orbits
We recall the weak amalgamation property (WAP), which will be crucial in the proof of Theorem 4.1. (WAP was first defined in [13], [11]. See [5, Section 2.4] for the formulation of WAP for strong classes.)
Let $\mathcal{D}$ be a reasonable class of $L^{+}$-expansions of an $L$-amalgamation class $(\mathcal{K},\leq)$. Recall that $(\mathcal{D},\leq)$ has the weak amalgamation property (WAP) if for all $A\in\mathcal{D}$, there exists $B\in\mathcal{D}$ and a $\leq$-strong $L^{+}$-embedding $f:A\to B$ such that, for any $\leq$-strong $L^{+}$-embeddings $f_{i}:B\to C_{i}\in\mathcal{D}$ ($i=0,1$), there exists $D\in\mathcal{D}$ and $\leq$-strong $L^{+}$-embeddings $g_{i}:C_{i}\to D$ ($i=0,1$) with $g_{0}\circ f_{0}\circ f=g_{1}\circ f_{1}\circ f$. (Note that here we specify only that the diagram commutes for $A$.)
Fact 2.3 ([5, Lemma 2.23]).
Let $\mathcal{D}$ be a reasonable class of $L^{+}$-expansions of an $L$-amalgamation class $(\mathcal{K},\leq)$ with Fraïssé limit $M$. Suppose that $X(\mathcal{D})$ is a minimal flow.
If $(\mathcal{D},\leq)$ does not have the weak amalgamation property, then all $\operatorname{Aut}(M)$-orbits on $X(\mathcal{D})$ are meagre.
For a proof, see [16, Lemma 1.77] (a straightforward correction of the proof in [5]).
@@ 3. Admissible orders: a Ramsey expansion of $(\mathcal{D}_{1},\sqsubseteq_{s})$
We now provide an explicit description of a Ramsey expansion of $(\mathcal{D}_{1},\sqsubseteq_{s})$, given by the admissible orders on $(\mathcal{D}_{1},\sqsubseteq_{s})$, using Theorem 1.4 of [6]. This will be an essential tool in the proof of 4.1. (This Ramsey expansion will also have the expansion property over $(\mathcal{D}_{1},\sqsubseteq_{s})$, though we will not use this.) In the below two definitions, we adapt definitions from [6, Section 1 and Section 3] to the specific case of $(\mathcal{D}_{1},\sqsubseteq_{s})$.
@@
(In the above, we adapt [6, Definition 3.5]. Several aspects of the general definition in [6] simplify in this case: (A4) can be omitted as closure components are single vertices, and (A6) follows from ()($\ast$)( ∗ ) and $(3)$.)
The below theorem is an immediate translation of [6, Theorem 1.4] to the context of this paper. We will explain how to adapt [6, Theorem 1.4] to our context at the end of this section.
@@
We now explain how to adapt [6, Theorem 1.4] and the definition of admissible orders found in [6] to give the definitions and theorem above. The paper [6] gives Ramsey expansions for classes of finite structures in languages that may include set-valued function symbols, which enables us to deal with the strong class $(\mathcal{D}_{1},\sqsubseteq_{s})$.
Definition 3.6 ([6, Section 1]).
@@
The above framework enables us to deal with $(\mathcal{D}_{1},\sqsubseteq_{s})$ as follows ([6, Section 5.1]). Let $\tilde{L}$ consist of the binary relational language $L_{\text{or}}$ of oriented graphs together with a unary set-valued function symbol $F$. Let $\tilde{\mathcal{D}}_{1}$ consist of the $\tilde{L}$ structures $\tilde{A}=(A,F_{A})$ where $A\in\mathcal{D}_{1}$ and $F_{A}:A\to\mathcal{P}(A)$ is a function sending each vertex of $A$ to its out-neighbourhood in $A$. Then there is a bijection $\mathcal{D}_{1}\to\tilde{\mathcal{D}}_{1}$ sending each $A\in\mathcal{D}_{1}$ to its unique $\tilde{L}$-expansion $\tilde{A}$ in $\tilde{\mathcal{D}}_{1}$, and for $A,B\in\mathcal{D}_{1}$, we have that $A\sqsubseteq_{s}B$ iff $\tilde{A}\subseteq\tilde{B}$. We then have that $\tilde{L}$-embeddings between elements of $\tilde{\mathcal{D}}_{1}$ are $\sqsubseteq_{s}$-embeddings when considered in the language $L_{\text{or}}$, and therefore $\tilde{\mathcal{D}}_{1}$ is a free amalgamation class.
We now recall [6, Theorem 1.4], which will give us an explicitly defined Ramsey expansion of $\tilde{\mathcal{D}}_{1}$ via admissibly ordered structures.
Theorem ([6, Theorem 1.4]).
@@
The above theorem, together with [6, Definition 3.5], which gives the explicit definition of admissible orders, gives a Ramsey expansion $\tilde{\mathcal{O}}_{1}$ of $\tilde{\mathcal{D}}_{1}$. Using the correspondence between $\tilde{\mathcal{D}}_{1}$ and $(\mathcal{D}_{1},\sqsubseteq_{s})$ detailed in the preceding paragraph, we thus obtain a Ramsey expansion $(\mathcal{O}_{1},\sqsubseteq_{s})$ of $(\mathcal{D}_{1},\sqsubseteq_{s})$ satisfying the conditions of 3.2 (this definition is just a direct adaptation of [6, Definition 3.5]).
@@ 4.1. Preparatory definitions and lemmas
Let $N_{1}=(M_{1},\rho)$ be the Fraïssé limit of $(\mathcal{D}_{1},\sqsubseteq_{s})$, and let $(N_{1},\prec_{\alpha})$ be the Fraïssé limit of $(\mathcal{O}_{1},\sqsubseteq_{s})$. (Here we use [5, Theorem 2.10]: $(\mathcal{D}_{1},\sqsubseteq_{s})$ is a strong expansion of $(\mathcal{C}_{1},\leq_{1})$ and $(\mathcal{O}_{1},\sqsubseteq_{s})$ is a strong expansion of $(\mathcal{D}_{1},\sqsubseteq_{s})$.)
Recall that $G_{1}=\operatorname{Aut}(M_{1})$. Let $H_{1}=\operatorname{Aut}(N_{1},\alpha)$. As $(\mathcal{O}_{1},\sqsubseteq_{s})$ has the Ramsey property by 3.3, by the fundamental result of the KPT correspondence ([12, Theorem 4.8]) formulated for strong classes ([5, Theorem 2.13]) we have that $H_{1}$ is extremely amenable.
@@ 4.2. Setup and proof notation
Let $Y\subseteq\mathcal{LO}(M_{1})$ be a minimal subflow of $G\curvearrowright\mathcal{LO}(M_{1})$. As $H$ is extremely amenable, the flow $H\curvearrowright Y$ has an $H$-fixed point $\prec_{\beta}$, and as $Y$ is a minimal $G$-flow, we have $Y=\overline{G\,\cdot\!\prec_{\beta}}$. Let $\mathcal{J}=\operatorname{Age}_{\leq_{1}}(M_{1},\prec_{\beta})$. By 2.1, we have $Y=X(\mathcal{J})$. We will show that $(\mathcal{J},\leq_{1})$ does not have the weak amalgamation property (WAP), which implies that all $G$-orbits on $Y$ are meagre by 2.3.